# Interpretable and flexible non-intrusive reduced-order models using reproducing kernel Hilbert spaces

### v1

@@ 4 Non-intrusive model reduction via kernel interpolation
where $\hat{\mathbf{q}}:[0,T]\to{}^{r}$ and $\hat{\mathbf{f}}:{}^{r}\to{}^{r}$. The structure of $\hat{\mathbf{f}}$ can be informed by intrusive projection, but unlike projection, defining $\hat{\mathbf{f}}$ through kernel interpolation does not require access to FOM operators such as $\mathbf{A}$ or $\mathbf{H}$ in eq. 3.14. We use the notation $\hat{\cdot}$ to mark non-intrusive objects and differentiate from intrusive objects, which are marked with $\tilde{\cdot}$.
@@ 4.1 Kernel reduced-order models
$\hat{\mathbf{q}}_{k}^{(\ell)}\coloneqq\mathbf{h}(\mathbf{q}_{k}^{(\ell)})=\mathbf{V}^{\mathsf{T}}(\mathbf{q}_{k}^{(\ell)}-\bar{\mathbf{q}})\in{}^{r}.$ (4.2)
@@
where $S$ is some set of functions and $R:S\to{}_{\geq 0}$ is a regularization function.
@@
where ${\boldsymbol{\Omega}}\in{}^{M(n_{t}+1)\times r}$ solves the linear system
@@
[…]
@@ 4.2 Specifying structure through kernel design
$\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)&=\mathbf{C}{\boldsymbol{\phi}}(\hat{\mathbf{q}}(t))\\ &=\hat{\mathbf{A}}\hat{\mathbf{q}}(t)+\hat{\mathbf{H}}_{2}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)]+\hat{\mathbf{H}}_{3}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)]+\hat{\mathbf{H}}_{4}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)],\end{aligned}$ (4.8)
@@
where $\tilde{\mathbf{b}}:{}^{n_{u}}\to{}^{r}$ and $\mathbf{u}:[0,T]\to{}^{n_{u}}$ model, for example, time-varying boundary conditions or forcing terms. In this case, we can construct feature maps ${\boldsymbol{\phi}}_{q}$ and ${\boldsymbol{\phi}}_{u}$ which aim to emulate the structures of $\tilde{\mathbf{f}}$ and $\tilde{\mathbf{b}}$, respectively, and define a concatenated feature map
@@
As discussed in Section 2.3, feature map kernels can lead to cost savings over generic kernels. Let $n_{\phi}$ be the dimension of the feature map, i.e., ${\boldsymbol{\phi}}(\hat{\mathbf{q}})\in{}^{n_{\phi}}$. Because the matrix $\mathbf{C}\in{}^{r\times n_{\phi}}$ can be computed once and reused, the cost of evaluating the ROM function $\hat{\mathbf{f}}$ online is $\mathcal{O}(rn_{\phi})$. Hence, if $n_{\phi}<Mn_{t}$, a feature map kernel is less expensive to evaluate than a generic kernel. If $n_{\phi}>Mn_{t}$ (e.g., due to a moderate reduced state dimension $r$), it can be beneficial to reduce $n_{\phi}$ and add a more generic element to the kernel to compensate. For instance, in place of the quartic ROM eq. 4.8, we may choose a quadratic feature map and add an RBF term to account for the cubic and quartic nonlinearities, resulting in a ROM of the form
@@
where ${\boldsymbol{\psi}}_{\!\epsilon}:{}^{r}\to{}^{M(n_{t}+1)}$ is as in eq. 2.11 and $c_{\psi}>0$ is a weighting coefficient as in eq. 2.15b. We test ROMs with this hybrid structure in Section 6. Note that this strategy can also apply to cases where the desired ROM structure is only partially known or representable by a feature map kernel.
@@ 4.3 Comparison to operator inference
for a specified feature map ${\boldsymbol{\phi}}:{}^{r}\to{}^{n_{\phi}}$ by solving the regularized residual minimization problem
$\min_{\hat{\mathbf{O}}\in{}^{r\times n_{\phi}}}\;\sum_{\ell=1}^{M}\sum_{k=0}^{n_{t}}\left\|\dot{\hat{\mathbf{q}}}_{k}^{(\ell)}-\hat{\mathbf{O}}{\boldsymbol{\phi}}(\hat{\mathbf{q}}_{k}^{(\ell)})\right\|_{2}^{2}+\|{\boldsymbol{\Gamma}}\hat{\mathbf{O}}^{\mathsf{T}}\|_{F}^{2},$ (4.14b)
where ${\boldsymbol{\Gamma}}\in{}^{n_{\phi}\times n_{\phi}}$. This is the generic learning problem eq. 4.4 with the function space $S$ given by
$S_{{\boldsymbol{\phi}}}=\left\{\mathbf{s}:{}^{r}\to{}^{r}\;:\;\mathbf{s}(\hat{\mathbf{q}})=\hat{\mathbf{O}}{\boldsymbol{\phi}}(\hat{\mathbf{q}})\quad\textup{for some}\quad\hat{\mathbf{O}}\in{}^{r\times n_{\phi}}\right\},$ (4.15)
@@
$\min_{\hat{\mathbf{A}}\in{}^{r\times r},\hat{\mathbf{H}}\in{}^{r\times r^{2}}}\;\sum_{\ell=1}^{M}\sum_{k=0}^{n_{t}}\left\|\dot{\hat{\mathbf{q}}}_{k}^{(\ell)}-\left(\hat{\mathbf{A}}\hat{\mathbf{q}}_{k}^{(\ell)}+\hat{\mathbf{H}}[\hat{\mathbf{q}}_{k}^{(\ell)}\otimes\hat{\mathbf{q}}_{k}^{(\ell)}]\right)\right\|_{2}^{2}+\left\|{\boldsymbol{\Gamma}}[~{}\hat{\mathbf{A}}~{}~{}\hat{\mathbf{H}}~{}]^{\mathsf{T}}\right\|_{F}^{2},$ (4.17)
and the solution $\hat{\mathbf{O}}=[~{}\hat{\mathbf{A}}~{}~{}\hat{\mathbf{H}}~{}]$ satisfies eq. 4.16 with $\mathbf{D}=[~{}\hat{\mathbf{Q}}^{\mathsf{T}}~{}~{}(\hat{\mathbf{Q}}\odot\hat{\mathbf{Q}})^{\mathsf{T}}~{}]\in{}^{M(n_{t}+1)\times(r+r^{2})}$. The underlying feature map in this case is the linear-quadratic map eq. 2.13. In practice, the compressed Kronecker product of Remark 3.1 is used so that eq. 4.17 has a unique solution.
@@
$\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)&=\hat{\mathbf{Z}}\big{(}K(\hat{\mathbf{Q}},\hat{\mathbf{Q}})+\gamma\mathbf{I}\big{)}^{-1}K(\hat{\mathbf{Q}},\hat{\mathbf{q}}(t))\\ &=\underbrace{\hat{\mathbf{Z}}\big{(}{\boldsymbol{\phi}}(\hat{\mathbf{Q}})^{\mathsf{T}}\mathbf{G}{\boldsymbol{\phi}}(\hat{\mathbf{Q}})+\gamma\mathbf{I}\big{)}^{-1}{\boldsymbol{\phi}}(\hat{\mathbf{Q}})^{\mathsf{T}}\mathbf{G}}_{\mathbf{C}}{\boldsymbol{\phi}}(\hat{\mathbf{q}}(t)),\end{aligned}$ (4.18)
@@ 5 Error estimates
We now derive several a posteriori error estimates for Kernel ROMs that relate the FOM solution $\mathbf{q}(t)$, the intrusive ROM solution $\tilde{\mathbf{q}}(t)$, and the Kernel ROM solution $\hat{\mathbf{q}}(t)$. These results require three main ingredients: the so-called local logarithmic Lipschitz constant, a Grönwall-type inequality, and standard error results for kernel interpolants. In this section, $\mathbf{M}\in{}^{n_{x}\times n_{x}}$ denotes a symmetric positive definite weighting matrix with Cholesky factorization $\mathbf{M}=\mathbf{L}\mathbf{L}^{\mathsf{T}}$. The $\mathbf{M}$-weighted inner product and norm are denoted with $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{\mathbf{M}}\coloneqq\mathbf{x}^{\mathsf{T}}\mathbf{M}\mathbf{y}=\left\langle\mathbf{L}^{\mathsf{T}}\mathbf{x},\mathbf{L}^{\mathsf{T}}\mathbf{y}\right\rangle$ and $\left\|\mathbf{x}\right\|_{\mathbf{M}}\coloneqq\left\langle\mathbf{x},\mathbf{x}\right\rangle_{\mathbf{M}}^{1/2}=\|\mathbf{L}^{\mathsf{T}}\mathbf{x}\|_{2}$, respectively.
@@ 5.1 Preliminaries
For a function $\mathbf{b}:{}^{n_{x}}\to{}^{n_{x}}$, the local logarithmic Lipschitz constant at $\mathbf{x}\in{}^{n_{x}}$ with respect to $\mathbf{M}$ is defined as
$\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})\coloneqq\sup_{\mathbf{z}\in{}^{d}}\frac{\left\langle\mathbf{z}-\mathbf{x},\mathbf{b}(\mathbf{z})-\mathbf{b}(\mathbf{x})\right\rangle_{\mathbf{M}}}{\left\|\mathbf{z}-\mathbf{x}\right\|_{\mathbf{M}}^{2}}.$ (5.1)
@@
The logarithmic norm of a matrix $\mathbf{B}\in{}^{n_{x}\times n_{x}}$ with respect to $\mathbf{M}$ is defined as
$\lambda_{\mathbf{M}}(\mathbf{B})\coloneqq\sup_{\mathbf{x}\in{}^{n_{x}}}\frac{\left\langle\mathbf{x},\mathbf{B}\mathbf{x}\right\rangle_{\mathbf{M}}}{\left\|\mathbf{x}\right\|_{\mathbf{M}}^{2}}=\max\sigma\left(\frac{1}{2}\left(\tilde{\mathbf{B}}+\tilde{\mathbf{B}}^{\mathsf{T}}\right)\right),$ (5.2)
@@
If $\mathbf{b}$ is an affine function, i.e., $\mathbf{b}(\mathbf{x})=\mathbf{B}\mathbf{x}+\mathbf{d}$ for some $\mathbf{B}\in{}^{n_{x}\times n_{x}}$ and $\mathbf{d}\in{}^{n_{x}}$, then $\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})=\lambda_{\mathbf{M}}(\mathbf{B})$. Note that the local logarithmic Lipschitz constant and the logarithmic norm can be negative, unlike a standard Lipschitz constant. We also note that if $\mathbf{b}$ is differentiable, then $\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})$ can be approximated by the logarithmic norm of the Jacobian $\mathbf{b}^{\prime}(\mathbf{x})$:
@@ 5.2 Error bounds
where $\mathbf{g}^{\prime}(\hat{\mathbf{q}})=\mathbf{V}+\mathbf{W}[\mathbf{I}\otimes\hat{\mathbf{q}}+\hat{\mathbf{q}}\otimes\mathbf{I}]\in{}^{n_{q}\times r}$ is the Jacobian of $\mathbf{g}$. Although we use a QM to define the reconstruction mapping $\mathbf{g}$, the following error analysis holds for any reconstruction mapping of the same structure, namely with $\mathbf{g}$ taken to be the sum of an affine part and a nonlinear part.
@@
[…]
@@
A caveat to the result in Theorem 5.1 is that it relies on Corollary 2.2, which requires zero regularization. However, as we demonstrate empirically in Section 6, the error bound eq. 5.5 still holds when the regularization hyperparameter $\gamma$ is small. Secondly, computing the local logarithmic Lipschitz constant $\lambda_{\mathbf{M}}[\mathbf{f}]$ is difficult to do in general. In practice, we instead approximate it using the logarithmic norm of $\mathbf{f}^{\prime}(\mathbf{g}(\hat{\mathbf{q}}(t)))$. Lastly, we note that the estimate eq. 5.5 requires evaluating the FOM right-hand side $\mathbf{f}$, and therefore is a code-intrusive error bound. We leave the non-intrusive estimation of the bound eq. 5.5 to future work.
We also obtain the following a posteriori error result for intrusive projection-based ROMs by examining the special case $\hat{\mathbf{f}}=\tilde{\mathbf{f}}$.
@@
The following error estimate holds for all $t\in[0,T]$:
[…]
We conclude with an error result comparing the intrusive projection-based ROM solution $\tilde{\mathbf{q}}(t)$ and the Kernel ROM solution $\hat{\mathbf{q}}(t)$. Let $\hat{\mathbf{e}}(t)=\tilde{\mathbf{q}}(t)-\hat{\mathbf{q}}(t)$, which satisfies the ODE
$\frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{e}}(t)=\tilde{\mathbf{f}}(\tilde{\mathbf{q}}(t))-\hat{\mathbf{f}}(\hat{\mathbf{q}}(t)),\qquad\hat{\mathbf{e}}(0)=\mathbf{0}.$ (5.8)
We then have the following.
Proposition 5.1.
Let $\hat{\mathbf{M}}\in{}^{r\times r}$ be a symmetric positive definite weighting matrix with Cholesky factorization $\hat{\mathbf{M}}=\hat{\mathbf{L}}\hat{\mathbf{L}}^{\mathsf{T}}$. If $\hat{\mathbf{f}}$ is an unregularized kernel interpolant of $\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\in{\cal H}_{K}^{r}$ where $\left\|{\boldsymbol{\delta}}(\hat{\mathbf{q}}(s))\right\|_{\hat{\mathbf{M}}}<\delta(s)$, then
[…]
Proof.
The dynamics in eq. 5.8 can be rewritten as
Therefore
Applying Lemma 5.1 yields the result. ∎

### v2

@@ 4 Non-intrusive model reduction via kernel interpolation
where $\hat{\mathbf{q}}:[0,T]\to\real^{r}$ and $\hat{\mathbf{f}}:\real^{r}\to\real^{r}$. The structure of $\hat{\mathbf{f}}$ can be informed by intrusive projection, but unlike projection, defining $\hat{\mathbf{f}}$ through kernel interpolation does not require access to FOM operators such as $\mathbf{A}$ or $\mathbf{H}$ in eq. 3.14. We use the notation $\hat{\cdot}$ to mark non-intrusive objects and differentiate from intrusive objects, which are marked with $\tilde{\cdot}$.
@@ 4.1 Kernel reduced-order models
$\hat{\mathbf{q}}_{k}^{(\ell)}\coloneqq\mathbf{h}(\mathbf{q}_{k}^{(\ell)})=\mathbf{V}^{\mathsf{T}}(\mathbf{q}_{k}^{(\ell)}-\bar{\mathbf{q}})\in\real^{r}.$ (4.2)
@@
where $S$ is some set of functions and $R:S\to\real_{\geq 0}$ is a regularization function.
@@
where ${\boldsymbol{\Omega}}\in\real^{M(n_{t}+1)\times r}$ solves the linear system
@@
$\begin{aligned} \hat{\mathbf{Q}}&=\begin{bmatrix}\hat{\mathbf{q}}_{0}^{(1)}&\cdots&\hat{\mathbf{q}}_{n_{t}}^{(1)}&\hat{\mathbf{q}}_{0}^{(2)}&\cdots&\hat{\mathbf{q}}_{n_{t}}^{(M)}\end{bmatrix}\in\real^{r\times M(n_{t}+1)},\\ \hat{\mathbf{Z}}&=\begin{bmatrix}\dot{\hat{\mathbf{q}}}_{0}^{(1)}&\cdots&\dot{\hat{\mathbf{q}}}_{n_{t}}^{(1)}&\dot{\hat{\mathbf{q}}}_{0}^{(2)}&\cdots&\dot{\hat{\mathbf{q}}}_{n_{t}}^{(M)}\end{bmatrix}\in\real^{r\times M(n_{t}+1)}.\end{aligned}$ (4.6c)
@@ 4.2 Specifying structure through kernel design
$\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)&=\mathbf{C}{\boldsymbol{\phi}}(\hat{\mathbf{q}}(t))\\ &=\hat{\mathbf{A}}\hat{\mathbf{q}}(t)+\hat{\mathbf{H}}_{2}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)]+\hat{\mathbf{H}}_{3}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)]+\hat{\mathbf{H}}_{4}[\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)\otimes\hat{\mathbf{q}}(t)],\end{aligned}$ (4.8)
@@
where $\tilde{\mathbf{b}}:\real^{n_{u}}\to\real^{r}$ and $\mathbf{u}:[0,T]\to\real^{n_{u}}$ model, for example, time-varying boundary conditions or forcing terms. In this case, we can construct feature maps ${\boldsymbol{\phi}}_{q}$ and ${\boldsymbol{\phi}}_{u}$ which aim to emulate the structures of $\tilde{\mathbf{f}}$ and $\tilde{\mathbf{b}}$, respectively, and define a concatenated feature map
@@
As discussed in Section 2.3, feature map kernels can lead to cost savings over generic kernels. Let $n_{\phi}$ be the dimension of the feature map, i.e., ${\boldsymbol{\phi}}(\hat{\mathbf{q}})\in\real^{n_{\phi}}$. Because the matrix $\mathbf{C}\in\real^{r\times n_{\phi}}$ can be computed once and reused, the cost of evaluating the ROM function $\hat{\mathbf{f}}$ online is $\mathcal{O}(rn_{\phi})$. Hence, if $n_{\phi}<Mn_{t}$, a feature map kernel is less expensive to evaluate than a generic kernel. If $n_{\phi}>Mn_{t}$ (e.g., due to a moderate reduced state dimension $r$), it can be beneficial to reduce $n_{\phi}$ and add a more generic element to the kernel to compensate. For instance, in place of the quartic ROM eq. 4.8, we may choose a quadratic feature map and add an RBF term to account for the cubic and quartic nonlinearities, resulting in a ROM of the form
@@
where ${\boldsymbol{\psi}}_{\!\epsilon}:\real^{r}\to\real^{M(n_{t}+1)}$ is as in eq. 2.11 and $c_{\psi}>0$ is a weighting coefficient as in eq. 2.15b. We test ROMs with this hybrid structure in Section 6. Note that this strategy can also apply to cases where the desired ROM structure is only partially known or representable by a feature map kernel.
@@ 4.3 Comparison to operator inference
for a specified feature map ${\boldsymbol{\phi}}:\real^{r}\to\real^{n_{\phi}}$ by solving the regularized residual minimization problem
$\min_{\hat{\mathbf{O}}\in\real^{r\times n_{\phi}}}\;\sum_{\ell=1}^{M}\sum_{k=0}^{n_{t}}\left\|\dot{\hat{\mathbf{q}}}_{k}^{(\ell)}-\hat{\mathbf{O}}{\boldsymbol{\phi}}(\hat{\mathbf{q}}_{k}^{(\ell)})\right\|_{2}^{2}+\|{\boldsymbol{\Gamma}}\hat{\mathbf{O}}^{\mathsf{T}}\|_{F}^{2},$ (4.14b)
where ${\boldsymbol{\Gamma}}\in\real^{n_{\phi}\times n_{\phi}}$. This is the generic learning problem eq. 4.4 with the function space $S$ given by
$S_{{\boldsymbol{\phi}}}=\left\{\mathbf{s}:\real^{r}\to\real^{r}\;:\;\mathbf{s}(\hat{\mathbf{q}})=\hat{\mathbf{O}}{\boldsymbol{\phi}}(\hat{\mathbf{q}})\quad\textup{for some}\quad\hat{\mathbf{O}}\in\real^{r\times n_{\phi}}\right\},$ (4.15)
@@
$\min_{\hat{\mathbf{A}}\in\real^{r\times r},\hat{\mathbf{H}}\in\real^{r\times r^{2}}}\;\sum_{\ell=1}^{M}\sum_{k=0}^{n_{t}}\left\|\dot{\hat{\mathbf{q}}}_{k}^{(\ell)}-\left(\hat{\mathbf{A}}\hat{\mathbf{q}}_{k}^{(\ell)}+\hat{\mathbf{H}}[\hat{\mathbf{q}}_{k}^{(\ell)}\otimes\hat{\mathbf{q}}_{k}^{(\ell)}]\right)\right\|_{2}^{2}+\left\|{\boldsymbol{\Gamma}}[~\hat{\mathbf{A}}~~\hat{\mathbf{H}}~]^{\mathsf{T}}\right\|_{F}^{2},$ (4.17)
and the solution $\hat{\mathbf{O}}=[~\hat{\mathbf{A}}~~\hat{\mathbf{H}}~$] satisfies eq. 4.16 with $\mathbf{D}=[~\hat{\mathbf{Q}}^{\mathsf{T}}~~(\hat{\mathbf{Q}}\odot\hat{\mathbf{Q}})^{\mathsf{T}}~]\in\real^{M(n_{t}+1)\times(r+r^{2})}$. The underlying feature map in this case is the linear-quadratic map eq. 2.13. In practice, the compressed Kronecker product of Remark 3.1 is used so that eq. 4.17 has a unique solution.
@@
$\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}\hat{\mathbf{q}}(t)&=\hat{\mathbf{Z}}\big(K(\hat{\mathbf{Q}},\hat{\mathbf{Q}})+\gamma\mathbf{I}\big)^{-1}K(\hat{\mathbf{Q}},\hat{\mathbf{q}}(t))\\ &=\underbrace{\hat{\mathbf{Z}}\big({\boldsymbol{\phi}}(\hat{\mathbf{Q}})^{\mathsf{T}}\mathbf{G}{\boldsymbol{\phi}}(\hat{\mathbf{Q}})+\gamma\mathbf{I}\big)^{-1}{\boldsymbol{\phi}}(\hat{\mathbf{Q}})^{\mathsf{T}}\mathbf{G}}_{\mathbf{C}}{\boldsymbol{\phi}}(\hat{\mathbf{q}}(t)),\end{aligned}$ (4.18)
@@ 5 Error estimates
We now derive several a posteriori error estimates for Kernel ROMs that relate the FOM solution $\mathbf{q}(t)$, the intrusive ROM solution $\tilde{\mathbf{q}}(t)$, and the Kernel ROM solution $\hat{\mathbf{q}}(t)$. Mathematical error estimates for non-intrusive ROMs are uncommon in the literature, typically restricted to special cases, and often difficult to obtain. Even in the intrusive case, error estimates for QM ROMs are not well-studied. One advantage of our kernel methods-based approach is that certain error estimates are straightforward to derive. This section presents an error bound for a general, non-intrusive kernel ROM with a generic nonlinear decoder $\mathbf{g}$. One immediate consequence is an intrusive error bound (Corollary 5.1), which we include to fill the gap in the literature for intrusive QM ROMs.
These results require three main ingredients: the so-called local logarithmic Lipschitz constant, a Grönwall-type inequality, and standard error results for kernel interpolants. In this section, $\mathbf{M}\in\real^{n_{x}\times n_{x}}$ denotes a symmetric positive definite weighting matrix with Cholesky factorization $\mathbf{M}=\mathbf{L}\mathbf{L}^{\mathsf{T}}$. The $\mathbf{M}$-weighted inner product and norm are denoted with $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{\mathbf{M}}\coloneqq\mathbf{x}^{\mathsf{T}}\mathbf{M}\mathbf{y}=\left\langle\mathbf{L}^{\mathsf{T}}\mathbf{x},\mathbf{L}^{\mathsf{T}}\mathbf{y}\right\rangle$ and $\left\|\mathbf{x}\right\|_{\mathbf{M}}\coloneqq\left\langle\mathbf{x},\mathbf{x}\right\rangle_{\mathbf{M}}^{1/2}=\|\mathbf{L}^{\mathsf{T}}\mathbf{x}\|_{2}$, respectively.
@@ 5.1 Preliminaries
For a function $\mathbf{b}:\real^{n_{x}}\to\real^{n_{x}}$, the local logarithmic Lipschitz constant at $\mathbf{x}\in\real^{n_{x}}$ with respect to $\mathbf{M}$ is defined as
$\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})\coloneqq\sup_{\mathbf{z}\in\real^{d}}\frac{\left\langle\mathbf{z}-\mathbf{x},\mathbf{b}(\mathbf{z})-\mathbf{b}(\mathbf{x})\right\rangle_{\mathbf{M}}}{\left\|\mathbf{z}-\mathbf{x}\right\|_{\mathbf{M}}^{2}}.$ (5.1)
@@
The logarithmic norm of a matrix $\mathbf{B}\in\real^{n_{x}\times n_{x}}$ with respect to $\mathbf{M}$ is defined as
$\lambda_{\mathbf{M}}(\mathbf{B})\coloneqq\sup_{\mathbf{x}\in\real^{n_{x}}}\frac{\left\langle\mathbf{x},\mathbf{B}\mathbf{x}\right\rangle_{\mathbf{M}}}{\left\|\mathbf{x}\right\|_{\mathbf{M}}^{2}}=\max\sigma\left(\frac{1}{2}\left(\tilde{\mathbf{B}}+\tilde{\mathbf{B}}^{\mathsf{T}}\right)\right),$ (5.2)
@@
If $\mathbf{b}$ is an affine function, i.e., $\mathbf{b}(\mathbf{x})=\mathbf{B}\mathbf{x}+\mathbf{d}$ for some $\mathbf{B}\in\real^{n_{x}\times n_{x}}$ and $\mathbf{d}\in\real^{n_{x}}$, then $\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})=\lambda_{\mathbf{M}}(\mathbf{B})$. Note that the local logarithmic Lipschitz constant and the logarithmic norm can be negative, unlike a standard Lipschitz constant. We also note that if $\mathbf{b}$ is differentiable, then $\Lambda_{\mathbf{M}}[\mathbf{b}](\mathbf{x})$ can be approximated by the logarithmic norm of the Jacobian $\mathbf{b}^{\prime}(\mathbf{x})$:
@@ 5.2 Error bounds
where $\mathbf{g}^{\prime}$ is the Jacobian of $\mathbf{g}$. Although we use a QM to define the reconstruction mapping $\mathbf{g}$, the following error analysis holds for any reconstruction mapping.
@@
$\left\|\mathbf{e}(t)\right\|_{\mathbf{M}}$ $\leq\int_{0}^{t}(\alpha_{P}(s)+\alpha_{K}(s))e^{\int_{s}^{t}\beta(\tau)d\tau}ds+e^{\int_{0}^{t}\beta(\tau)d\tau}\left\|\mathbf{e}_{N}(0)\right\|_{\mathbf{M}},$ $\forall$ $\;t\in[0,T],$ (5.5)
@@
A caveat to the result in Theorem 5.1 is that it relies on Corollary 2.2, which requires zero regularization. However, as we demonstrate empirically in Section 6, the error bound eq. 5.5 still holds when the regularization hyperparameter $\gamma$ is small. Secondly, computing the local logarithmic Lipschitz constant $\lambda_{\mathbf{M}}[\mathbf{f}]$ is difficult to do in general. In practice, we instead approximate it using the logarithmic norm of $\mathbf{f}^{\prime}(\mathbf{g}(\hat{\mathbf{q}}(t)))$. Similarly, since $\tilde{\mathbf{f}}$ is unavailable, we estimate $\|\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\|_{{\cal H}_{K}^{r}}$ with $\|\hat{\mathbf{f}}\|_{{\cal H}_{K}^{r}}$, which is easily computable. Lastly, we note that the estimate eq. 5.5 requires evaluating the FOM right-hand side $\mathbf{f}$ and its Jacobian. While these evaluations are possible with some black-box PDE solvers, they are not possible in general, which limits the utility of Theorem 5.1 for the a posteriori certification of a Kernel ROM. Nonetheless, the error estimate eq. 5.5 provides mathematical justification for the usage of Kernel ROMs, and our numerical experiments empirically verify the estimate’s validity. Using Theorem 5.1 for ROM certification would require the non-intrusive estimation of eq. 5.5, i.e., assuming that $\mathbf{f}$ and $\mathbf{f}^{\prime}$ cannot be queried, which is non-trivial and thus left to future work.
We conclude with an a posteriori error result for intrusive projection-based ROMs by examining the special case $\hat{\mathbf{f}}=\tilde{\mathbf{f}}$. As mentioned above, error estimates for intrusive QM ROMs are not widely available in the literature, so we provide the following result, which includes intrusive QM ROMs as the special case when $\mathbf{g}$ is of the form eq. 3.5.
@@
If $\hat{\mathbf{f}}=\tilde{\mathbf{f}}$, then the following error estimate holds for all $t\in[0,T]$:
$\left\|\mathbf{q}(t)-\mathbf{g}(\tilde{\mathbf{q}}(t))\right\|_{\mathbf{M}}$ $\leq\int_{0}^{t}\left\|(\mathbf{I}-\mathbf{g}^{\prime}(\tilde{\mathbf{q}}(s))\mathbf{V}^{\mathsf{T}})\mathbf{f}(\mathbf{g}(\tilde{\mathbf{q}}(s)))\right\|_{\mathbf{M}}e^{\int_{s}^{t}\Lambda_{\mathbf{M}}[\mathbf{f}](\mathbf{g}(\hat{\mathbf{q}}_{I}(\tau)))d\tau}ds$ (5.7)
$\quad+e^{\int_{0}^{t}\Lambda_{\mathbf{M}}[\mathbf{f}](\mathbf{g}(\tilde{\mathbf{q}}(\tau)))d\tau}\left\|\mathbf{q}(0)-\mathbf{g}(\hat{\mathbf{q}}(0))\right\|_{\mathbf{M}}.$